The sun is a sphere with an estimated mass of 1.80×1030 kg.

If the radius of the sun is 7.001×105 km, what is the average density of the sun in units of grams per cubic centimeter? The volume of a sphere is (4/3)π r3.
Answer #1

Sol :-

Mass of the sun (m) = 1.80 × 1030 kg = 1.80 × 1033 g

Radius of the sun (r) = 7.001 × 105 km = 7.001 × 1010 cm

Volume of the sun (V) = (4/3)π.r3

= (4/3) × (3.14) × (7.001 × 1010 cm)3

= 1.4366 × 1033 cm3

Now, Average density of the sun in g/cm3 = Mass of the sun (m) / Volume of the sun (V)

= 1.80 × 1033 g / 1.4366 × 1033 cm3

= 1.25 g/cm3

Hence, Average density of the sun = 1.25 g/cm3